The Gilster Company, a machine tooling firm, has several plants. One plant, located in St. Falls, Minnesota, uses a job order costing system for its batch production processes. The St. Falls plant has two departments through which most jobs pass. Plantwide overhead, which includes the plant manager’s salary, accounting personnel, cafeteria, and human resources, is budgeted at $250,000. During the past year, actual plantwide overhead was $240,000. Each department’s overhead consists primarily of depreciation and other machine-related expenses. Selected budgeted and actual data from the St. Falls plant for the past year are as follows.

Department A Department B
Budgeted department overhead
(excludes plantwide overhead) $ 150,000 $ 600,000
Actual department overhead 160,000 620,000
Expected total activity:
Direct labor hours 35,000 15,000
Machine-hours 10,000 40,000
Actual activity:
Direct labor hours 51,000 9,000
Machine-hours 10,500 42,000

For the coming year, the accountants at St. Falls are in the process of helping the sales force create bids for several jobs. Projected data pertaining only to job no. 110 are as follows.

Direct materials $ 25,000
Direct labor cost:
Department A (2,200 hr) 45,000
Department B (800 hr) 10,000
Machine-hours projected:
Department A 200
Department B 1,200
Units produced 10,000

Required:

a-1. Assume the St. Falls plant uses a single plantwide overhead rate to assign all overhead (plantwide and department) costs to jobs. Find the overhead rate by using expected direct labor hours.

a-2. Determine the projected amount of total manufacturing costs per unit for the units in job no. 110.
Answer #1
a-1) Overhead rate hour = total budgeted overhead/budgeted labor hours
total budgeted overhead (250000+150000+600000) = 1000000
budgeted labor hour (35000+15000) = 50000
Overhead rate hour = 1000000/50000 =20 per direct labor hour
a-2) total manufacturing cost per unit job 110
Direct material 25000
direct labor
Department A 45000
Department B 10000
overhead (2200+800)*20 60000
total cost 140000
unit produced 10000
manufacturing cost per unit 14

Please give a thumbs up if it is helpful & let me know if any doubt